OpenStudy (korosh23):

Math question! Factor using bum method 3xsquare+5x+2 Please, factor completely. Thank you.

OpenStudy (jhannybean):

bum method??

OpenStudy (jhannybean):

Can you describe this method to me?Haha

OpenStudy (mathmate):

For BUM method, see: http://www.bcmath.ca/m9h/m9hch4/m9pch4.4/index.html (Have to click on the slide to show each sentence)

OpenStudy (jhannybean):

Oh, that's what that method is called? Haha

OpenStudy (jhannybean):

\[\begin{align} \color{red}{3}x^2+5x+\color{red}{2} \\ &= x^2+5x+6 \\ \\ &= (x+3)(x+2) \\ \\ &=\left(x+\frac{3}{3}\right)\left(x+\frac{2}{3}\right) \\ \\ & \sf \text{divide by the leading coefficient to keep the equation constant} \\ &\sf \text{simplify whicever fractions you can, and for the ones you cannot} \\ &\sf\text{ multiply diagonally to the variable.} \\ \\ &=\boxed{(x+1)(3x+2)} \end{align}\]

OpenStudy (korosh23):

Thank you.